# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

### v2

@@ 3.3. Interviews
In addition to the large-scale analysis of Q&A posts, we conducted semi-structured interviews to gain a deeper understanding of practitioners’ day-to-day experiences with Ansible. In the following section, we describe the recruitment process, the structure of the interviews, and the analysis. All interview materials are in the Supplemental Materials.
3.3.1. Participants
We categorize participants based on their level of expertise with Ansible. For this we use years of experience due to support in previous work (LaToza and Myers, 2010; Roehm et al., 2012) and also to mitigate the Dunning-Kruger Effect, where users overestimate their skills in self-assessments (Mahmood, 2016). We define three levels of expertise with Ansible: Beginner (less than two years of experience, n = 5); Intermediate (two to five years of experience, inclusive, n = 7); and Experienced (more than five years of experience, n = 8).
@@ 3.3.1. Participants
We recruited 20 participants through a pre-screening survey distributed via email and various social media platforms, including LinkedIn, X (formerly Twitter), Discord, and Slack ( 1 in Figure 2). To meet our eligibility criteria, participants had to have prior experience using Ansible. Following this stage, participants selected their preferred interview modality: a text-based or video-based session conducted via Zoom. The availability of both formats was intended to address known challenges in recruiting software developers for research studies (Baltes and Diehl, 2016), a challenge we also encountered in our study. A total of 118 out of 127 survey participants were eligible and, among these, 20 were available for an interview (an interview rate of 15.7%). Most of the participants (12 out of 20) opted for a text interview. Several participants mentioned that they would only be able to participate in the interview if they could do it through text. As an incentive, the interviewed participants could enter a raffle for a \$100 gift card.
The supplementary material contains full details on the demographics of the participants. In summary, participants reported anywhere from 1 to more than 9 years of Ansible experience, with many currently using Ansible (15). Most identified as male (17), ranged in age from 18–34 (11), and held at least a bachelor’s degree (17). Almost all had a job in CS and had a formal CS education (17).
@@ 3.3.2. Interviews
We conducted semi-structured interviews to obtain a nuanced understanding of the participants’ experiences with Ansible ( 2 in Figure 2). Semi-structured interviews are the most widely used interviewing method (DiCicco-Bloom and Crabtree, 2006) and allow adaptability to explore emerging themes and seek clarifications (Lazar et al., 2017). We piloted the protocol with four participants. Both text and voice-based interviews required around 30 minutes of attention time. The latter were recorded and automatically transcribed with participant consent using Zoom’s built-in transcription functionality. Research team members manually reviewed and corrected the resulting transcripts. Each interview followed a structured protocol that encompassed five key thematic areas: contextual background, perceived advantages, challenges, recommendations, and future of Ansible. The interview started with general background questions. We then moved the discussion towards the advantages of Ansible and adoption challenges. Then, we focused the discussion on Ansible’s limitations. The interview concluded with exploring recommendations to improve Ansible and future perspectives. We actively sought clarification whenever responses were ambiguous or required elaboration, both in textual and voice-based interviews.
3.3.3. Analysis
@@
We developed five emergent codebooks using thematic analysis (Braun and Clarke, 2006; Lazar et al., 2017), following a bottom-up approach that allowed themes and codes to emerge inductively from our data. We started by reading the transcripts to highlight recurring ideas, creating initial codes, and iteratively refining them into coherent themes ( 3 and 4 in Figure 2). Our codebooks are related to the five key thematic areas as follows: (1) Background: Exploring why participants use Ansible and the specific contexts in which they apply it; (2) Advantages: Identifying the key motivations behind adoption; (3) Challenges: Documenting difficulties and limitations; (4) Recommendations: Capturing participants’ wishes; (5) Future: Addressing potential future developments.
@@
For each of these five areas, we developed a codebook. We followed a methodology similar to the one present in prior work (Jackson et al., 2024; Stalnaker et al., 2024; Karlsson et al., 2007), and that is recommended by Human-Centered research guidelines (Lazar et al., 2017). One of the authors developed an initial codebook that captured key aspects of the responses. The initial codebook was shared with a second coder (also an author) , who reviewed and suggested modifications based on a subset of interview data. We repeated this process multiple times until consensus was reached and the codebook was deemed stable. Once we established the final version of the codebook, we double-coded all of the interviews, where one coder systematically applied the codes to all interview responses, and the second coder independently reviewed the applied codes. This review involved cross-checking all coded transcripts to identify inconsistencies or overlooked themes. The second coder made detailed notes on any disagreements. When there was a change to the codebook, for example, a new code, the coding process was restarted. Finally, both coders met to discuss discrepancies and reach a consensus ( 5 and 6 in Figure 2). When attributing quotes, we report the participant ID and the experience of the participant: (B)eginner, (I)ntermediate or (E)xperienced. For example, B1 refers to a beginner participant.
@@ 4. RQ1: Common Issues in Ansible Usage
Figure 3. Distribution of posts across topics and their respective themes.
@@
This section presents the seven primary challenge categories encountered by Ansible practitioners, obtained from the topic modeling analysis and interview data, in descending order of prevalence.
4.1. Control/Data Flow and Execution
According to our topic modeling analysis, Control/Data Flow and Execution is the most significant theme that represents the challenges Ansible practitioners face. Our approach identified 8,793 posts related to Control Flow and Execution, which represents around 37.9% of all recent posts.
4.1.1. Topic Analysis
@@
Within the Control/Data Flow and Execution theme, Loops represent the most significant challenge for Ansible practitioners (16.4%). As one practitioner stated , “I don’t understand why loops are so hard to do in this language.”. Loops are followed by SSH Connection Issues to hosts (14.9%), Variable Management (14.4%), Conditional Execution (12.4%), and Undefined Variables (12.3%).
Topic Evolution
The number of references to Loops in posts increased from 2.9% of all posts in 2013 to 7.4% in 2022, before decreasing to 4.3% in 2024 ( ). Contrarily, the number of posts related to Variable Management decreased from 8.9% in 2013 to 5.4% in 2024 ( ). This evolution suggests that, while loop issues persist and even increase, Ansible may be making progress in addressing the challenges related to management of variables.
4.1.2. Interviews
@@
Participants described various control flow issues. For instance, B1 mentioned frustration with sequential execution slowing runs and a lack of synchronization primitives (n = 2). In more advanced setups, complexities grew around managing dependencies (n = 3). Beyond these granular concerns, a handful of participants reflected on code structure/organization difficulties (n = 5). In contrast, others cited Ansible’s inherently stateless (n = 5) nature as a double-edged sword — simpler to run but lacking built-in features for preserving or comparing states. The push model/agentless (n = 5) approach was both praised and criticized: praised for minimal overhead yet lamented for complicating continuous enforcement.
4.2. Reliability
Reliability emerged as the second most discussed theme, accounting for approximately 32.3% of all recent posts.
4.2.1. Topic Analysis
Although Error Handling is the most represented topic within the Reliability theme (16.4%), TopicGPT’s interpretation of this topic is quite generic. As a result, despite its description being related to the handling of errors in Ansible playbooks, the topic was frequently used in cases where arbitrary errors occurred. Following Error Handling, the most discussed issues are related to file management (11.5%), permissions (8.1%), JMESPath (5.9%) and the use of shell commands (5.4%).
Topic Evolution
In 2013, Permission Issues represented just 1.4% of the total posts in that year, but by 2024, this figure had increased to 3.2% ( ). This increase may suggest that Ansible practitioners are encountering more permission-related issues, likely due to interactions with hosts that have more strict security configurations.
4.2.2. Interviews
Reliability issues emerged in discussions around debugging. Multiple participants reported that Ansible is hard to debug (n = 10). I7 tries to avoid Ansible to debug complex cases: “if there’s some more complex stages in my playbooks what I try to do is to transfer those to a shell script that I can pass outside of Ansible and I can test those separately.”. B2 described moments when “it would get stuck for 15 minutes in a supposedly 1 min script before I forced it to close without any visible changes on the terminal.”. Some lamented the lack of infrastructure debugging tools (n = 5) built into Ansible, while others pointed to bad error messages when modules returned cryptic outputs. As B4 put it, “I actually believe that there’s not a proper way to debugging this [Infrastructure as Code] type of scripts and tools.”.
4.3. Lack of Support
@@
Lack of Support is the second most significant theme, accounting for 32.1% of the recent posts with a total of 7,443 posts.
4.3.1. Topic Analysis
@@
Within this theme, AWX-related issues dominated discussions (20.9%), followed by module execution failures (12.7%), integration issues (10.8%), module usage problems (9.8%), and Windows-related challenges (9.1%). These findings suggest that practitioners encounter persistent challenges with the use of AWX, an open-source project serving as the upstream for Ansible’s enterprise version, along with issues related to module reliability, understanding how to use modules, integrating Ansible playbooks with external systems, and managing Windows hosts.
Topic Evolution
The number of posts related to AWX, launched in 2017, has increased significantly from 1.6% of the overall posts in its release year to 9.8% in 2024 ( ). Another topic that saw significant growth was Integration Issues, increasing from 0.8% in 2013 to 5.1% in 2024 ( ). This suggests that as Ansible matured, practitioners increasingly sought solutions for integrating with external services. In the opposite direction, posts related to Module Development decreased from 1.6% in 2013 to 0.6% in 2024 ( ).
4.3.2. Interviews
@@
The theme lack of support touched on problems such as missing features that require modules (n = 3), limited OS support (n = 4), and Python version constraints (n = 1). Other participants criticized the lack of support for certain modules (n = 7). As E4 put it “There’s a lack of support for some modules (…). You often have to write your own or hope someone did it.”.
4.4. Usability
Usability represents approximately 16.6% of recent posts, totaling 3,846 posts.
4.4.1. Topic Analysis
@@
The issues related to Jinja2 are by far the most represented, with around 33.1% of the posts within the Usability theme. For instance, in one of the posts, an Ansible practitioner mentioned: “Unfortunately, Jinja2 has been incredibly challenging to work with.”. Issues related to quoting (22.4%), playbook inputs (20.4%), YAML syntax (10.8%), and regex (8.2%) are also prominently discussed. When talking about their issues with YAML, a practioner mentioned: “I find YAML or Ansible’s interpretation of Yaml very frustrating and confusing.”.
Topic Evolution
In 2013, issues related to Jinja2 were not as prominent, only 2.4% of the posts that year. However, Jinja2-related issues peaked at 6.9% in 2021, before decreasing to 3.8% by 2024 ( ).
4.4.2. Interviews
@@
Ansible’s usability emerged as a significant issue for many participants. It includes user onboarding issues and complaints about complex syntax. Participants expressed that it can be challenging for beginners (n = 12) to grasp the tool’s declarative nature, in particular for those that lacked a coding background (n = 2). Others felt they had to make a mental shift (n = 4), especially if they were used to more imperative scripting. As E8 said, “If you’ve never seen Terraform, CloudFormation (…) it’ll look weird and alienating.”. Moreover, sometimes it can be hard to pick the right approach (n = 4). Participants also complained about complex syntax in general (n = 4). Ansible’s reliance on YAML and Jinja introduced the problems of YAML and Jinja complexity (n = 5 and 6). As E5 said, “I once spent 8 hours of work figuring out the escaping on a line of SQL because YAML is so bad at figuring out quotes (…) The hell that is escaping in YAML will never die.”.
4.5. Performance and Scalability
@@
Performance and Scalability represents approximately 7.5% of recent posts, with 1,738 posts discussing these issues.
4.5.1. Topic Analysis
@@
Issues related to dynamic inventories (31.0%) and inventory management (29.8%) are the most frequently discussed topics within Performance and Scalability, followed by performance (17.7%), remote execution (8.5%) and delegation (7.7%).
Topic Evolution
Inventory management issues have remained relatively stable between 2013 and 2024, consistently accounting for between 1.8% and 2.5% of the total yearly posts ( ).
4.5.2. Interviews
@@
Scalability limitations (7) were frequently mentioned by participants managing large infrastructures. B1 stated: “When you run a playbook for like 100 hosts, it gets really confusing trying to make sense of the messages.”. Participants specifically described problems related to managing inventories (n = 1) and code awareness (n = 4) in large projects. A related source of concern was slow execution (n = 8). While some liked the simplicity of SSH-based workflows, a few individuals cited agentless (n = 4) as detrimental in high-volume scenarios. E4 was particularly direct: “It does not scale when you start to manage 100th of servers, it becomes slow because there is no agent.”.
4.6. Maintenance
@@
Maintenance challenges appear in 683 posts, which corresponds to about 3% of the total of recent posts.
4.6.1. Topic Analysis
@@
Challenges related to deprecated features and settings in Ansible (34.4%), dependency management (23.3%), compatibility issues (19.6%) — which include, for example, supporting legacy operating systems —, different Ansible versions (14.2%) or version controlling Ansible projects (12.3%) are the most relevant topics in the Maintenance theme.
4.6.2. Interviews
@@
Participants noted that certain aspects of Ansible maintenance remain nontrivial. In particular, several mentioned that removing configurations is difficult (n = 3) — as modifying the playbook alone does not guarantee that older changes get undone without explicit tasks. E2 explained that “Doing things like removing config from hosts tends not to work so well in Ansible,”. Configuration drift (n = 5) was also mentioned. As I3 described “Deletes are hard: one cannot merely delete parts of the Ansible playbook and expect those changes to be undone in hosts. Instead, we must change or even add tasks undoing the configuration we want gone.”. Other participants mentioned that Ansible can be hard to maintain (n = 5) and were concerned about deprecations (n = 5).
4.7. Information Gaps
@@
The Information Gaps theme is the least prevalent, accounting for only 465 posts, or about 2% of all recent posts. Of these 465 posts, 464 address the challenges practitioners face with documentation.
4.7.1. Interviews
@@
Another concern revolved around information gaps, particularly documentation that is either overwhelming or incomplete. E2 talked about misleading information: “Another issue is incorrect or misleading answers on forums and newsgroups. Beginners may struggle to distinguish good advice from bad.”. The theme of documentation gaps (n = 9) appeared in several responses, with participants describing the difficulty of locating relevant or clear information among Ansible’s extensive (but sometimes inconsistent) references. Although multiple participants praised Ansible’s core documentation (described in Section 5), they also identified a steep learning curve for interpreting specific advanced.

### v3

@@ 3.3. Interviews
In addition to the large-scale analysis of Q&A posts, we conducted semi-structured interviews to gain deeper insight into practitioners’ experiences with Ansible.
Participants
Participants were categorized by Ansible expertise using years of experience, following prior work (LaToza and Myers, 2010; Roehm et al., 2012), and to mitigate self-assessment bias, such as the Dunning–Kruger effect (Mahmood, 2016). We define three levels: Beginner (<2<2 years, n=5), Intermediate ($2\geq$ and $\leq 5$ years, n=7), and Experienced (>5>5 years, n=8).
@@
We recruited 20 participants through a pre-screening survey distributed via email and various social media platforms, including LinkedIn, X (formerly Twitter), Discord, and Slack ( 1 in Fig. 2). To meet our eligibility criteria, participants had to have prior experience using Ansible. Following this stage, participants selected their preferred interview modality: a text-based or video-based session conducted via Zoom. The availability of both formats was intended to address known challenges in recruiting software developers for research studies (Baltes and Diehl, 2016), a challenge we also encountered in our study. A total of 118 of 127 survey participants were eligible, and among them, 20 were available for an interview (an interview rate of 15.7%). Most of the participants (12 out of 20) opted for a text interview. Several participants mentioned that they would only be able to participate in the interview if it were conducted via text. As an incentive, the interviewed participants could enter a raffle for a \$100 gift card. Demographic details are provided in the Supplemental Materials. In summary, participants reported 1 to over 9 years of Ansible experience; most were active users (15), identified as male (17), were aged 18–34 (11), held at least a bachelor’s degree (17), and had CS education and employment (17).
@@
We conducted semi-structured interviews to obtain a nuanced understanding of the participants’ experiences with Ansible ( 2 in Fig. 2). Semi-structured interviews are the most widely used interviewing method (DiCicco-Bloom and Crabtree, 2006) and allow adaptability to explore emerging themes and seek clarifications (Lazar et al., 2017). We piloted the protocol with four participants. Both text and voice-based interviews required around 30 minutes of attention time. The latter were recorded and automatically transcribed with participant consent using Zoom’s built-in transcription functionality. Research team members manually reviewed and corrected the resulting transcripts. Each interview followed a structured protocol that encompassed four key thematic areas: perceived advantages, challenges, recommendations, and future of Ansible. The interview started with general background questions. We then moved the discussion towards the advantages of Ansible and adoption challenges. Then, we focused the discussion on Ansible’s limitations. The interview concluded with exploring recommendations to improve Ansible and future perspectives. We actively sought clarification whenever responses were ambiguous or required elaboration, both in textual and voice-based interviews.
Analysis
@@
We developed four emergent codebooks using thematic analysis (Braun and Clarke, 2006; Lazar et al., 2017), following a bottom-up approach that allowed themes and codes to emerge inductively from our data. We started by reading the transcripts to highlight recurring ideas, creating initial codes, and iteratively refining them into coherent themes ( 3 and 4 in Fig. 2). Our codebooks are related to the four key thematic areas as follows: (1) Advantages: Identifying the key motivations behind adoption; (2) Challenges: Documenting difficulties and limitations; (3) Recommendations: Capturing participants’ wishes; (4) Future: Addressing potential future developments.
@@
For each of these four areas, we developed a codebook. We followed a methodology similar to the one present in prior work (Jackson et al., 2024; Stalnaker et al., 2024; Karlsson et al., 2007), and that is recommended by Human-Centered research guidelines (Lazar et al., 2017). One of the authors developed an initial codebook that captured key aspects of the responses. The initial codebook was shared with a second coder (also an author), who reviewed and suggested modifications based on a subset of interview data. We repeated this process multiple times until consensus was reached and the codebook was deemed stable. Once we established the final version of the codebook, we double-coded all of the interviews, where one coder systematically applied the codes to all interview responses, and the second coder independently reviewed the applied codes. This review involved cross-checking all coded transcripts to identify inconsistencies or overlooked themes. The second coder made detailed notes on any disagreements. When there was a change to the codebook, for example, a new code, the coding process was restarted. Finally, both coders met to discuss discrepancies and reach a consensus ( 5 and 6 in Fig. 2). When attributing quotes, we report the participant ID and the experience of the participant: (B)eginner, (I)ntermediate or (E)xperienced. For example, B1 refers to a beginner participant. See Supplemental Materials for the full codebooks.
@@ 4. RQ1: Common Issues in Ansible Usage
Figure 3. Topic distribution (Feb 2020–Feb 2025). Sparklines show yearly percentage trends across the full Ansible post history.
@@
This section presents the seven primary challenge categories encountered by Ansible practitioners, ordered by prevalence. Figure 3 shows the distribution of posts by topic and theme.
Control/Data Flow and Execution
is the most prevalent challenge category for Ansible practitioners. This category accounts for 10,721 posts, representing 44.4% of recent discussions.
Topic Analysis
@@
Loops represent the most significant challenge for Ansible practitioners (13.9%). As one practitioner stated in the posts, “I don’t understand why loops are so hard to do in this language.” Loops are followed by SSH Connection Issues to hosts (12.7%), Variable Management (12.3%), Conditional Execution (10.6%), and Undefined Variables (10.4%).
Interviews
@@
Participants described various control flow issues. For instance, B1 mentioned frustration with sequential execution slowing runs and a lack of synchronization primitives (n=2). In more advanced setups, complexities grew around managing dependencies (n=3). Beyond these granular concerns, a handful of participants reflected on code structure/organization difficulties (n=5). In contrast, others cited Ansible’s inherently stateless (n=5) nature as making it simpler to run but also lacking built-in features for preserving or comparing states. The push model/agentless (n=5) was both praised for minimal overhead and criticized for complicating continuous enforcement.
Lack of Support
@@
is the second most significant theme, accounting for 38.2% of the recent posts with a total of 9,226 posts.
Topic Analysis
@@
AWX-related issues, where AWX is the open-source upstream project for Ansible Tower, dominated discussions (17.3%), followed by module execution failures (10.7%), integration issues (8.5%), module usage problems (8.2%), and Windows-related challenges (7.7%).
Interviews
@@
Interviewees reported missing features that require custom modules (n=3) and limited OS support (n=4). Other participants criticized the lack of support for certain modules (n=7). As E4 put it “There’s a lack of support for some modules (…). You often have to write your own or hope someone did it.”
Reliability
emerged as the third most discussed theme, accounting for approximately 37.2% of all recent posts.
Topic Analysis
Although Error Handling is the most represented topic (14.2%), TopicGPT’s interpretation of it is generic. As a result, despite its description being related to the handling of errors in Ansible playbooks, the topic was frequently used in cases where arbitrary errors occurred. Following Error Handling, the most discussed issues are related to file management (10.0%), permissions (7.0%), JMESPath (5.1%) and the use of shell commands (4.8%).
Interviews
Reliability issues primarily emerged around debugging. Multiple participants reported that Ansible is hard to debug (n=10). I7 tries to avoid Ansible to debug complex cases: “if there’s some more complex stages in my playbooks what I try to do is to transfer those to a shell script that I can pass outside of Ansible and I can test those separately.” B2 described moments when “it would get stuck for 15 minutes in a supposedly 1 min script before I forced it to close without any visible changes on the terminal.” Some lamented the lack of built-in infrastructure debugging tools (n=5), while others pointed to bad error messages (n=5) when modules returned cryptic outputs. As B4 put it, “I actually believe that there’s not a proper way to debugging this type of scripts and tools.”
Usability
represents ~18.8% of recent posts (3,846 posts).
Topic Analysis
@@
Issues related to Jinja2 are the most prevalent, accounting for 29.1% of usability-related posts. Issues related to quoting (19.7%), playbook inputs (17.9%), YAML syntax (9.8%), and regex (7.2%) are also prominently discussed. When talking about their issues with YAML, a practitioner mentioned: “I find YAML or Ansible’s interpretation of YAML very frustrating and confusing.”
Interviews
@@
Ansible’s usability emerged as a significant concern for many participants. Participants described Ansible as challenging for beginners (n=12), particularly for those without a coding background (n=2). Others felt they had to make a mental shift (n=4), especially if they were used to more imperative scripting. As E8 said, “If you’ve never seen Terraform, CloudFormation (…) it’ll look weird and alienating”. Some participants found it hard to pick the right approach (n=4). Participants also complained about complex syntax (n=4). Ansible’s reliance on YAML and Jinja introduced the problems of YAML and Jinja complexity (n=5 and 6). As E5 said, “I once spent 8 hours of work figuring out the escaping on a line of SQL (…) The hell that is escaping in YAML will never die.”
Performance and Scalability
@@
represents approximately 8.1% of recent posts, with 1,965 posts discussing these issues.
Topic Analysis
@@
Dynamic inventories (29.0%) and inventory management (27.3%) dominate discussions, followed by performance (16.3%), remote execution (7.7%), and delegation (7.0%).
Interviews
@@
Participants managing large infrastructures frequently reported scalability limitations (n=7). B1 stated: “When you run a playbook for like 100 hosts, it gets really confusing trying to make sense of the messages”. Participants specifically described problems related to code awareness (n=4) in large projects. A related source of concern was slow execution (n=8). While some liked the simplicity of SSH-based workflows, a few individuals cited agentless (n=4) as detrimental in high-volume scenarios. E4 was particularly direct: “It does not scale when you start to manage 100th of servers, it becomes slow because there is no agent.”
Maintenance
@@
challenges appear in 750 posts, which corresponds to about 3.1% of the total of recent posts.
Topic Analysis
@@
Challenges related to deprecated features and settings in Ansible (32.7%), dependency management (22.3%), compatibility issues (18.5%) (e.g., with legacy OSs), different Ansible versions (13.5%) or version controlling Ansible projects (11.5%) are the most relevant topics in the Maintenance theme.
Interviews
@@
Participants noted that removing configurations is difficult (n=3) as modifying the playbook alone does not guarantee that older changes get undone. E2 explained that “Doing things like removing config from hosts tends not to work so well in Ansible,”. Configuration drift (n=5) was also mentioned. As I3 described “Deletes are hard: one cannot merely delete parts of the Ansible playbook and expect those changes to be undone in hosts. Instead, we must change or even add tasks undoing the configuration we want gone.” Other participants mentioned that Ansible can be hard to maintain (n=5) and were concerned about deprecations (n=5).
Information Gaps
@@
is the least prevalent theme, accounting for only 482 posts, or about 2% of all recent posts. Of these 482 posts, 480 address the challenges practitioners face with documentation.
Interviews
@@
Another concern revolved around information gaps, particularly incomplete documentation. E2 talked about misleading information: “Another issue is incorrect or misleading answers on forums and newsgroups. Beginners may struggle to distinguish good advice from bad.” Participants that talked about documentation gaps (n=9) described the difficulty of locating relevant or clear information among Ansible’s extensive (but sometimes inconsistent) references. Although multiple participants praised Ansible’s core documentation (described in Section 5), they also identified a steep learning curve for interpreting advanced instructions.
4.1. Topic Evolution
Figure 3 shows the yearly percentage evolution of each topic over the entire Ansible post history until 2024. Most topics exhibit a stable prevalence over time. However, issues related to AWX, integration and compatibility issues, and documentation increase in prominence, whereas topics such as variable management and module development show a gradual decrease.